In a collge there are 13 times as many students as professors. If together the students and professors number 28,000. how many students are there in the college?
Answer by checkley77(12844)

X+13X=28,000
14X=28,000
X=28,000/14
X=2,000 NUMBER OF PROFESSORS.
13*2,000=26,000 NUMBER OF STUDENTS.
PROOF:
2,000+26,000=28,000
28,000=28,000